# arXiv:2603.03264v1[astro-ph.SR] 03 Mar 2026

### 03 Mar 2026 v1

@@ Abstract
Context. Observational constraints on stellar magnetic fields are essential to both stellar and planetary physics: they play a key role in the formation and evolution of stars and planets, and are responsible for spurious signals in radial velocity curves that impact the detection and characterization of exoplanets. Recent observations have revealed the diversity and evolution of large-scale magnetic fields in low-mass stars. These large-scale fields, however, only account for a small fraction of the observed unsigned magnetic flux. The other crucial stellar magnetism information originate from the small (spatial) scale magnetic fields, which accounts for most of the surface magnetic flux and exhibits clear temporal evolution of time scales of years.
@@
Aims. With this work, we aim at developing new fast techniques to extract small-scale magnetic field estimates from time series of observed high-resolution spectra. One objective is to develop tools that will allow the community to take full advantage of the upcoming monitoring surveys carried out with various high-resolution spectrometers. Our ultimate goal is to study the temporal evolution of small-scale magnetic fields which will provide insight into the magnetic properties of low-mass stars and their magnetic cycles.
@@
Methods. We implement a process to capture relative pixel variations caused by changes in magnetic field strengths, relying on synthetic spectra computed with ZeeTurbo. This approach provides extremely fast and reliable estimates of relative magnetic field strength variations from series of high-resolution spectra, mitigating the impact of systematics between models and observations. We assess the performance of the proposed method through its application to simulated data and publicly available observed spectra recorded with SPIRou, Narval and ESPaDOnS. In addition, we implement a model-driven process to derive relative temperature variations and explore the influence magnetic fields have on these measurements.
@@
Results. Our results are in excellent agreement with magnetic field estimates previously obtained from spectra recorded with SPIRou. The method provides robust constraints on the structure of the magnetic field variations, and proves to be relatively insensitive to small changes in the assumed atmospheric parameters and broadening. We find that magnetic field variations have the potential of introducing biases in relative temperature estimates, in particular for the Narval/ESPaDOnS spectral domains, containing a large number of magnetically-sensitive transitions and where contrast is more important. Our application to archival data provides new constraints on the evolution of small-scale magnetic fields, and underscores the potential of the proposed method for the analysis of datasets in the context of large observation programs.
@@
Conclusions. By reducing the problem to a set of linear equations, our method offers extremely fast results, making it viable for integration in future pipelines developed for large spectroscopic surveys. These estimates will provide much needed information to correct radial velocity curves and constrain dynamo process.
@@ 1 Introduction
Sun-like stars and M dwarfs are known to host magnetic fields (e.g., Saar & Linsky, 1985; Johns-Krull & Valenti, 1996; Kochukhov, 2021; Reiners et al., 2022), with consequences on their polarized and unpolarized spectra. These fields are thought to be responsible for activity phenomena, introducing spurious signal in radial velocity (RV) measurements (Dumusque et al., 2021; Bellotti et al., 2022) and consequently impacting exoplanet detection and characterization. Magnetic fields also play a crucial role in stellar formation and evolution processes (Donati & Landstreet, 2009), leading to angular momentum loss (e.g., Skumanich, 1972; Vidotto et al., 2014) and surface inhomogeneities (spots, plages and faculae).
@@
M dwarfs are the most numerous stars in the solar neighborhood (Reylé et al., 2021), and are considered prime targets for the search of habitable exoplanets. The precise and accurate characterization of these stars is essential to extract reliable estimates on the masses and radii of the exoplanets orbiting them (Bonfils et al., 2013; Dressing & Charbonneau, 2015; Gaidos et al., 2016). The masses of M dwarfs range from 0.08 to 0.57 $\rm M_{\odot}$ (Pecaut & Mamajek, 2013): stars with masses below 0.35 $\rm M_{\odot}$ are believed to be fully convective (Chabrier & Baraffe, 1997), while M dwarfs with masses above 0.35 $\rm M_{\odot}$ have radiative interiors and an outer convective envelope separated by a transition region known as the tachocline (Spiegel & Zahn, 1992). The tachocline is thought to play a key role in the 22-year magnetic cycle observed in the Sun (see Parker, 1955; Charbonneau, 2010). In fully convective M dwarfs, dynamo processes must be revisited and alternative scenarios have been proposed to generate large-scale fields without the need of a tachocline (Chabrier & Küker, 2006; Yadav et al., 2015).
@@
The last decade has been marked by the rapid development of high-resolution spectrometers, including several near-infrared instruments such as CARMENES (Quirrenbach et al., 2014), CRIRES+ (Dorn et al., 2023) and SPIRou (Donati et al., 2020), motivated in part by the search for exoplanets around M dwarfs. The spectropolarimetric capabilities of some of these instruments led to numerous studies focused on large-scale stellar magnetic fields (e.g., Kochukhov & Reiners, 2020; Finociety et al., 2023; Donati et al., 2023b; Bellotti et al., 2024), while unpolarized spectra have allowed the community to characterize the small-scale magnetic fields of low-mass stars (e.g., Shulyak et al., 2017; Reiners et al., 2022; Cristofari et al., 2023a, b). Combined efforts to characterize both large-scale and small-scale magnetic fields have provided a complete picture of cool stars magnetism (e.g., Kochukhov & Lavail, 2017; Donati et al., 2023a, 2025; Bellotti et al., 2025). Recent works have reported efforts to characterize the variation of small-scale magnetic fields throughout years-long observation campaigns (Donati et al., 2023a; Cristofari et al., 2025b), revealing rotational modulation and long-term fluctuations. Small-scale magnetic fields have been proposed as an excellent proxy to correct for activity jitters in radial velocity curves (Haywood et al., 2016, 2022). The extraction of small-scale field fluctuations from the high-resolution spectra, and in particular from the spectra used to obtain radial velocity curves, is therefore particularly timely.
@@
With this work, we propose a new process to derive variations of the average surface magnetic field strength ($\langle B\rangle$) throughout time series of spectra. We validate our method and assess its performance on simulated spectra before applying it to archival observed spectra. Section 2 describes the observational data used in this study. Our method and its implementation are described in sections 3 and 4. The application of our method to observed spectra is described in section 5, and our results are discussed in section 6.
@@ 2 Observations and data reduction
This work takes advantage of public high-resolution spectroscopic observations obtained with three instruments: ESPaDOnS (Donati et al., 2006) and SPIRou (Donati et al., 2020) installed at the Canada-France Hawaii-Telescope (CFHT), and Narval installed on the Télescope Bernard Lyot (TBL) at the Pic du Midi observatory, France (Donati, 2003).
ESPaDOnS and Narval data were reduced with the LIBRE-ESPRIT reduction software (Donati et al., 1997), and the continuum normalized spectra were retrieved from Polarbase (Petit et al., 2014). SPIRou data were reduced with APERO (Cook et al., 2022) and downloaded from the Canadian Astrophysics Data Center (CADC). APERO provides multiple data products including spectra before and after correction of telluric lines (with extention e.fits and t.fits, respectively) along with blaze profiles estimated from flat-field exposures. An estimate of the photon noise was obtained from the number counts prior to telluric correction, and used to estimate uncertainties in the telluric-corrected spectra.
For this work, we refined our initial spectrum normalization procedure of SPIRou spectra used by Cristofari et al. (2023b, 2025a). For each observation, the spectral orders were corrected by the blaze profiles, and a 1D spectrum was obtained by merging all orders. Where spectral orders overlapped, only the segments with the highest throughput were kept to avoid biases of the normalization by noisy order edges. The blaze corrected spectrum was then divided in windows of equal velocity width (which we empirically fixed to 3000 $\rm km\,s^{-1}$). In each window, we computed the flux 95th percentile, and a cubic interpolation between the obtained points provides us with the continuum of the whole spectrum. Our process successfully captures low-frequency variation in the spectra, while preserving the structure of the spectral lines. We find that this process is particularly useful to preserve the shape of spectral lines recorded on an order edge, which can bias normalization in order-per-order normalization schemes.
Here we consider 3 M dwarfs: EV Lac, DS Leo and Barnard’s star. All their available SPIRou spectra were retrieved, and combined for each observation night to achieve higher signal to noise ratios. Most of the observations recorded with SPIRou were obtained in the context of the SPIRou Legacy Survey. For EV Lac and DS Leo, we additionally retrieved all of the available Narval and ESPaDOnS data. Regions with deep telluric absorption bands were excluded from the analysis. Because both instruments have similar characteristics, observations were used together, and spectra were combined each night to obtain higher signal to noise ratios. For EV Lac, this yielded spectra for 67 nights between 2005 to 2016, while we obtained spectra for 61 nights for DS Leo between 2006 and 2014.
@@ 3 Model description
The idea of extracting information about the stellar surface from relative variations in spectral features has been around for decades, with Gray & Johanson (1991); Gray (1994) selecting highly $T_{\rm eff}$-sensitive line ratios to derive temperature variations, and recent studies focused on deriving relative radial velocity variations to achieve the meter-per-second precisions needed to detect and characterize exoplanets (see, e.g., Artigau et al., 2022; Shahaf & Zackay, 2023). More recently, Artigau et al. (2024) proposed to derive relative temperature variations from a fully-empirical approach relying on the formalism adopted by Bouchy et al. (2001) and Artigau et al. (2022). We begin by briefly reviewing the basis of this formalism.
We consider a spectrum A\@vec{A}, function of a parameter $x$ which is assumed to have a single value for all pixels (e.g. effective temperature, metallicity, magnetic field intensity). Small variations between A\@vec{A} and a reference spectrum $\@vec{A}_{\rm ref}$ can be approximated by the derivative of A\@vec{A} with respect to $x$ evaluated at $x_{\rm ref}$, so that:
@@
In the context of magnetic fields, however, the implementation of this method leads to spurious results, which we attribute to two main factors. First, for most pixels, variations in flux as a function of magnetic field in the 0–10 kG range are poorly modeled by a low-order polynomial, and the assumption of small variations fails. Secondly, observed spectra are poorly modeled by a single magnetic field intensity (see, e.g., Kochukhov, 2021; Reiners et al., 2022; Cristofari et al., 2023a, b). To overcome these limitations, one needs to adopt a more complete formalism describing the variation of pixels with magnetic field intensity for different magnetic components.
@@ 3.1 Formalism
In this section, we adopt the now commonly used formalism (see, e.g. Shulyak et al., 2014, 2019; Reiners et al., 2022; Cristofari et al., 2023a, b), in which a spectrum A\@vec{A} is represented as:
$\@vec{A}=\sum{a_{k}\@vec{F}_{k}}$ (2)
where $\@vec{F}_{k}$ is the synthetic disk-integrated spectrum of a star with an assumed radial magnetic field strength, and $a_{k}$ is the filling factor associated with that component. Previous studies suggest a range of useful components for M dwarfs spanning 0 kG to $12$ kG (Shulyak et al., 2019; Reiners et al., 2022; Cristofari et al., 2023a). Because the filling factors must add up to one for the problem to remain physical, this relation can be written as:
$\@vec{A}=\@vec{F}_{0}+\sum{a_{k}(\@vec{F}_{k}-\@vec{F}_{0})}$ (3)
where $\@vec{F}_{0}$ is the synthetic disk-integrated spectrum computed for a 0 kG magnetic field. The synthetic spectrum A\@vec{A} therefore depends solely on the filling factors $a_{k}$ associated with each magnetic component. Let us consider a reference spectrum, $\@vec{A}_{\rm ref}$, described by the filling factors $\epsilon_{k}$, so that $\@vec{A}_{\rm ref}=\@vec{F}_{0}+\sum{\epsilon_{k}(\@vec{F}_{k}-\@vec{F}_{0})}$}. Then a variation between A\@vec{A} and $\@vec{A}_{\rm ref}$ can be written:
$\@vec{A}-\@vec{A}_{\rm ref}=\sum{(a_{k}-\epsilon_{k})(\@vec{F}_{k}-\@vec{F}_{0})}$ (4)
The variation in flux can be described by a variation in filling factors $a_{k}-\epsilon_{k}$, which we will write $\delta a_{k}$ in the rest of the paper. The variation in surface averaged magnetic field strength $\langle B\rangle$ is then simply obtained as $\delta\langle B\rangle=\sum\delta a_{k}B_{k}$, where $B_{k}$ is the magnetic field strength for the $k$-th component.
Solving the equation 4 allows one to derive $\delta a_{k}$, provided that $A$, $A_{\rm ref}$ and $F_{k}$ for each magnetic components are known. In our application to observed spectra, the reference spectrum $A_{\rm ref}$ is computed from the observed spectra directly, so that the left hand-side of equation 4 comes from observations, while the right-hand side of the equation is informed by models (see Sec. 4).
@@ 4 Implementation and simulations
Our implementation relies on synthetic spectra computed with ZeeTurbo (Cristofari et al., 2023a) to obtain $\@vec{F}_{k}-\@vec{F}_{0}$ for different field strengths. Synthetic spectra were computed with higher sampling than observations, so that for an observed spectrum A\@vec{A}, one may choose the pixels in $\@vec{F}_{k}-\@vec{F}_{0}$ corresponding to the nearest observed wavelengths. We build the reference spectrum $\@vec{A}_{\rm ref}$ by taking the median of the observed spectra in the stellar reference frame, which we call ”template” in the rest of the paper.
Our process is inherently sensitive to spurious pixels in the observations, and flux fluctuations not properly accounted for by error bars (such as telluric line residuals). To mitigate the impact of spurious pixels, we quadratically add the standard deviation of the flux across the series of observed spectra to the photon noise associated with each pixel. We then implement an iterative process to reject bad pixels. Specifically, the weighted Least-squares problem (equation 4) is solved using all points to obtain a first estimate of the $\delta a_{k}$, which is used to make a prediction on the variation of each pixel’s flux. We then compute the reduced $\chi^{2}$ on the predicted and observed flux variations, and use this value to inflate the error bars associated with each pixel. Each pixel for which the flux deviates by more than 3 times the inflated error bar is rejected, and the weighted least-square problem is solved on the remaining pixels. The process is repeated until all spurious pixels are rejected. Convergence is typically reached within the first 10 iterations in our applications. This approach provides a robust and conservative way to reject bad pixels that could impact the results. We find that this method is particularly robust to provide estimates for spectra with significant artifacts.

### 26 Mar 2026 v2

# A fast method for deriving relative small-scale magnetic field variations from high-resolution spectroscopy

@@ Abstract
Context. Setting observational constraints on stellar magnetic fields is essential for both stellar and planetary physics. They play a key role in the formation and evolution of stars and planets, and they are responsible for spurious signals in radial velocity curves that impact the detection and characterization of exoplanets. Recent observations have revealed the diversity and evolution of large-scale magnetic fields in low-mass stars. However, these large-scale fields only account for a small fraction of the observed unsigned magnetic flux. The other crucial stellar magnetism information originates from (spatially) small-scale magnetic fields, which account for most of the surface magnetic flux and exhibit a clear temporal evolution on timescales of many years.
@@
Aims. With this work, we aim to develop new fast techniques to extract small-scale magnetic field estimates from time series of observed high-resolution spectra. One objective is to develop tools that will enable the community to take full advantage of the upcoming monitoring surveys carried out with various high-resolution spectrometers. Our ultimate goal is to study the temporal evolution of small-scale magnetic fields and offer insights into the magnetic properties of low-mass stars and their magnetic cycles.
@@
Methods. We implemented a process to capture relative pixel variations caused by changes in magnetic field strengths, relying on synthetic spectra computed with ZeeTurbo. This approach provides extremely fast and reliable estimates of relative magnetic field strength variations from series of high-resolution spectra, mitigating the impact of systematics between models and observations. We assessed the performance of the proposed method through its application to simulated data and publicly available observed spectra recorded with SPIRou, Narval, and ESPaDOnS. In addition, we implemented a model-driven process to derive relative temperature variations and we explored the influence magnetic fields have on these measurements.
@@
Results. Our results are in excellent agreement with the magnetic field estimates previously obtained from spectra recorded with SPIRou. This method provides robust constraints on the structure of the magnetic field variations and proves to be relatively insensitive to small changes in the assumed atmospheric parameters and broadening. We find that magnetic field variations have the potential of introducing biases in relative temperature estimates. This is particularly relevant in the case of the Narval/ESPaDOnS spectral domains, which contain a large number of magnetically sensitive transitions and where contrast is more important. Our application to archival data provides new constraints on the evolution of small-scale magnetic fields and underscores the potential of the proposed method for analyzing data in the context of large observation programs.
@@
Conclusions. By reducing the problem to a set of linear equations, our method offers extremely fast results, making it viable for integration in future pipelines developed for large spectroscopic surveys. These estimates will provide much needed information to correct radial velocity curves and constrain dynamo processes.
@@ 1 Introduction
Sun-like stars and M dwarfs are known to host magnetic fields (e.g., Saar and Linsky, 1985; Johns-Krull and Valenti, 1996; Kochukhov, 2021; Reiners et al., 2022) that have consequences on their polarized and unpolarized spectra. These fields are thought to be responsible for activity phenomena, introducing spurious signal in radial velocity (RV) measurements (Dumusque et al., 2021; Bellotti et al., 2022) and consequently impacting exoplanet detection and characterization. Magnetic fields also play a crucial role in stellar formation and evolution processes (Donati and Landstreet, 2009), leading to angular momentum loss (e.g., Skumanich, 1972; Vidotto et al., 2014) and surface inhomogeneities (e.g., spots, plages, and faculae).
@@
M dwarfs are the most numerous stars in the solar neighborhood (Reylé et al., 2021) and they are considered prime targets for the search of habitable exoplanets. The precise and accurate characterization of these stars is essential to extract reliable estimates on the masses and radii of the exoplanets orbiting them (Bonfils et al., 2013; Dressing and Charbonneau, 2015; Gaidos et al., 2016). The masses of M dwarfs range from 0.08 to 0.57 $\rm M_{\odot}$ (Pecaut and Mamajek, 2013): stars with masses below 0.35 $\rm M_{\odot}$ are believed to be fully convective (Chabrier and Baraffe, 1997), while M dwarfs with masses above 0.35 $\rm M_{\odot}$ have radiative interiors and an outer convective envelope separated by a transition region known as the tachocline (Spiegel and Zahn, 1992). The tachocline is thought to play a key role in the 22-year magnetic cycle observed in the Sun (see Parker, 1955; Charbonneau, 2010). In fully convective M dwarfs, dynamo processes must be revisited and alternative scenarios have been proposed to generate large-scale fields without the need for a tachocline (Chabrier and Küker, 2006; Yadav et al., 2015).
@@
The last decade has been marked by the rapid development of high-resolution spectrometers, including several near-infrared instruments such as CARMENES (Quirrenbach et al., 2014), CRIRES+ (Dorn et al., 2023), and SPIRou (Donati et al., 2020), motivated in part by the search for exoplanets around M dwarfs. The spectropolarimetric capabilities of some of these instruments have led to numerous studies focused on large-scale stellar magnetic fields (e.g., Kochukhov and Reiners, 2020; Finociety et al., 2023; Donati et al., 2023b; Bellotti et al., 2024), while unpolarized spectra have enabled the community to characterize the small-scale magnetic fields of low-mass stars (e.g., Shulyak et al., 2017; Reiners et al., 2022; Cristofari et al., 2023a, b). Combined efforts to characterize both large-scale and small-scale magnetic fields have provided a complete picture of cool stars magnetism (e.g., Kochukhov and Lavail, 2017; Donati et al., 2023a, 2025; Bellotti et al., 2025). Recent works have reported efforts to characterize the variation of small-scale magnetic fields throughout observation campaigns lasting several years (Donati et al., 2023a; Cristofari et al., 2025) and revealing rotational modulation and long-term fluctuations. Small-scale magnetic fields have been proposed as an excellent proxy to correct for activity jitters in radial velocity curves (Haywood et al., 2016, 2022). The extraction of small-scale field fluctuations from the high-resolution spectra, particularly from the spectra used to obtain radial velocity curves, is therefore timely.
@@
With this work, we propose a new process for deriving variations in the average surface magnetic field strength ($\langle B\rangle$) throughout time series of spectra. We validate our method and assess its performance on simulated spectra before applying it to archival observed spectra. Section 2 describes the observational data used in this study. Our method and its implementation are described in Sects. 3 and 4. The application of our method to observed spectra is described in Sect. 5 and our results are discussed in Sect. 6.
@@ 2 Observations and data reduction
This work takes advantage of public high-resolution spectroscopic observations obtained with three instruments: ESPaDOnS (Donati et al., 2006) and SPIRou (Donati et al., 2020), installed at the Canada-France Hawaii-Telescope (CFHT), and Narval installed on the Télescope Bernard Lyot (TBL) at the Pic du Midi observatory, France (Donati, 2003).
ESPaDOnS and Narval data were reduced with the LIBRE-ESPRIT reduction software (Donati et al., 1997) and the continuum-normalized spectra were retrieved from Polarbase (Petit et al., 2014). SPIRou data were reduced with APERO (Cook et al., 2022) and downloaded from the Canadian Astrophysics Data Center (CADC). APERO provides multiple data products including spectra before and after correction of telluric lines (with extention e.fits and t.fits, respectively), along with blaze profiles estimated from flat-field exposures. An estimate of the photon noise was obtained from the number counts prior to telluric correction and used to estimate uncertainties in the telluric-corrected spectra.
For this work, we refined our initial spectrum normalization procedure of SPIRou spectra, used by Cristofari et al. (2023b, 2025). For each observation, the spectral orders were corrected by the blaze profiles and a 1D spectrum was obtained by merging all orders. In cases where spectral orders overlapped, only the segments with the highest throughput were kept to avoid biases of the normalization by noisy order edges. The blaze-corrected spectrum was then divided in windows of equal velocity width (which we empirically fixed to 3000 $\rm km\,s^{-1}$). In each window, we computed the flux 95th percentile and a cubic interpolation between the obtained points provided us with the continuum of the whole spectrum. Our process successfully captured low-frequency variation in the spectra, while preserving the structure of the spectral lines. We find that this process is particularly useful in preserving the shape of spectral lines recorded on an order edge, which can bias normalization in order-per-order normalization schemes.
Here, we consider three M dwarfs: EV Lac, DS Leo, and Barnard’s star. All their available SPIRou spectra were retrieved and combined for each observation night to achieve higher signal-to-noise ratios (S/Ns). Most of the observations recorded with SPIRou were obtained in the context of the SPIRou Legacy Survey. For EV Lac and DS Leo, we additionally retrieved all of the available Narval and ESPaDOnS data. Regions with deep telluric absorption bands were excluded from the analysis. Because both instruments have similar characteristics, observations were used together, and spectra were combined each night to obtain higher S/Ns. For EV Lac, this yielded spectra for 67 nights between 2005 to 2016 and for DS Leo, this yielded spectra for 61 nights between 2006 and 2014.
@@ 3 Model description
The idea of extracting information about the stellar surface from relative variations in spectral features has been around for decades, with Gray and Johanson (1991); Gray (1994) selecting highly $T_{\rm eff}$-sensitive line ratios to derive temperature variations. Some recent studies have been focused on deriving relative radial velocity variations to achieve the meter-per-second precisions needed to detect and characterize exoplanets (see, e.g., Artigau et al., 2022; Shahaf and Zackay, 2023). More recently, Artigau et al. (2024) proposed deriving relative temperature variations from a fully data-driven approach and relying on the formalism adopted by Bouchy et al. (2001) and Artigau et al. (2022). We begin by briefly reviewing the basis of this formalism.
We consider a spectrum, A\@vec{A}, function of a parameter, $x$, which is assumed to have a single value for all pixels (e.g., effective temperature, metallicity, magnetic field intensity). Small variations between A\@vec{A} and a reference spectrum, $\@vec{A}_{\rm ref}$, can be approximated by the derivative of A\@vec{A} with respect to $x$ evaluated at $x_{\rm ref}$, so that
@@
In the context of magnetic fields, however, the implementation of this method leads to spurious results, which we attribute to two main factors. First, for most pixels, variations in flux as a function of magnetic field in the 0–10 kG range are poorly modeled by a low-order polynomial; as a result, the assumption of small variations fails. Secondly, observed spectra are poorly modeled by a single magnetic field intensity (see, e.g., Kochukhov, 2021; Reiners et al., 2022; Cristofari et al., 2023a, b). To overcome these limitations, it is necessary to adopt a more complete formalism describing the variation of pixels with magnetic field intensity for different magnetic components.
@@ 3.1 Formalism
In this section, we describe the now commonly used formalism (see, e.g. Shulyak et al., 2014, 2019; Reiners et al., 2022; Cristofari et al., 2023a, b) that we adopted. Here, spectrum A\@vec{A} is represented as
$\@vec{A}=\sum{a_{k}\@vec{F}_{k}},$ (2)
where $\@vec{F}_{k}$ is the synthetic disk-integrated spectrum of a star with an assumed radial magnetic field strength and $a_{k}$ is the filling factor associated with that component. Previous studies suggest a range of useful components for M dwarfs spanning 0 kG to $12$ kG (Shulyak et al., 2019; Reiners et al., 2022; Cristofari et al., 2023a). Because the filling factors must add up to 1 for the problem to remain physical, this relation can be written as
$\@vec{A}=\@vec{F}_{0}+\sum{a_{k}(\@vec{F}_{k}-\@vec{F}_{0})},$ (3)
where $\@vec{F}_{0}$ is the synthetic disk-integrated spectrum computed for a 0 kG magnetic field. The synthetic spectrum A\@vec{A} therefore depends solely on the filling factors, $a_{k}$, associated with each magnetic component. Let us consider a reference spectrum, $\@vec{A}_{\rm ref}$, described by the filling factors, $\epsilon_{k}$, so that $\@vec{A}_{\rm ref}=\@vec{F}_{0}+\sum{\epsilon_{k}(\@vec{F}_{k}-\@vec{F}_{0})}$}. Then a variation between A\@vec{A} and $\@vec{A}_{\rm ref}$ can be written as
$\@vec{A}-\@vec{A}_{\rm ref}=\sum{(a_{k}-\epsilon_{k})(\@vec{F}_{k}-\@vec{F}_{0})}.$ (4)
The variation in flux can be described by a variation in filling factors $a_{k}-\epsilon_{k}$, which we denote as $\delta a_{k}$ hereafter. The variation in surface averaged magnetic field strength, $\langle B\rangle$, can then simply be obtained as $\delta\langle B\rangle=\sum\delta a_{k}B_{k}$, where $B_{k}$ is the magnetic field strength for the $k$-th component.
Solving Eq. 4 allows us to derive $\delta a_{k}$, provided that the values of $A$, $A_{\rm ref}$, and $F_{k}$ for each magnetic components are known. In our application to observed spectra, the reference spectrum, $A_{\rm ref}$, was computed from the observed spectra directly, so the left hand-side of Eq. 4 comes from observations, while the right-hand side of the equation is informed by models (see Sect. 4).
@@ 4 Implementation and simulations
Our implementation relies on synthetic spectra computed with ZeeTurbo (Cristofari et al., 2023a), which allowed us to obtain $\@vec{F}_{k}-\@vec{F}_{0}$ for different field strengths. Synthetic spectra were computed with higher sampling than observations, so that for an observed spectrum A\@vec{A}, we can choose the pixels in $\@vec{F}_{k}-\@vec{F}_{0}$ corresponding to the nearest observed wavelengths. We built the reference spectrum $\@vec{A}_{\rm ref}$ by taking the median of the observed spectra in the stellar reference frame, which we refer to as the template hereafter. Our process is inherently sensitive to spurious pixels in the observations and flux fluctuations not properly accounted for by error bars (e.g., telluric line residuals). To mitigate the impact of spurious pixels, we quadratically added the standard deviation of the flux across the series of observed spectra to the photon noise associated with each pixel. We then implemented an iterative process to reject bad pixels. Specifically, the weighted least-squares problem (Eq. 4) can be solved using all points to obtain a first estimate of the $\delta a_{k}$, which is used to make a prediction on the variation of each pixel’s flux. We then computed the reduced $\chi^{2}$ on the predicted and observed flux variations and used this value to inflate the error bars associated with each pixel. We rejected each pixel for which the flux ends up deviating by more than three times the inflated error bar and the weighted least-square problem was solved on the remaining pixels. The process was repeated until all spurious pixels are rejected. Convergence was typically reached within the first ten iterations in our applications. This approach provides a robust and conservative way to reject bad pixels that could impact the results. We find that this method is particularly robust to provide estimates for spectra with significant artifacts.